A quiz cosists of 20 multiple-choice questions, each with 6 possible answers. For someone who makes random guesses for all of the answers, find the probability of passing if the minimum passing grade is 40 %.
Answer by stanbon(75887)

A quiz cosists of 20 multiple-choice questions, each with 6 possible answers. For someone who makes random guesses for all of the answers, find the probability of passing if the minimum passing grade is 40 %
----
Binomial Problem with n = 20 and p(correct) = 1/6
---
20% of 20 = 4
----
P(4 <= x <= 20)
= 1 - P(0 <= x <= 3)
= 1-0.5665
= 0.4335
============
Cheers,
Stan H.
============